# VideoAds for Fast-Paced Video Understanding: Where Opensource Foundation Models Beat GPT-4o & Gemini-1.5 Pro

### v1

@@ Abstract
Advertisement videos serve as a rich and valuable source of purpose-driven information, encompassing high-quality visual, textual, and contextual cues designed to engage viewers. They are often more complex than general videos of similar duration due to their structured narratives and rapid scene transitions, posing significant challenges to multi-modal large language models (MLLMs). In this work, we introduce VideoAds, the first dataset tailored for benchmarking the performance of MLLMs on advertisement videos. VideoAds comprises well-curated advertisement videos with complex temporal structures, accompanied by manually annotated diverse questions across three core tasks: visual finding, video summary, and visual reasoning. We propose a quantitative measure to compare VideoAds against existing benchmarks in terms of video complexity. Through extensive experiments, we find that Qwen2.5-VL-72B, an opensource MLLM, achieves 73.35% accuracy on VideoAds, outperforming GPT-4o (66.82%) and Gemini-1.5 Pro (69.66%); the two proprietary models especially fall behind the opensource model in video summarization and reasoning, but perform the best in visual finding. Notably, human experts easily achieve a remarkable accuracy of 94.27%. These results underscore the necessity of advancing MLLMs’ temporal modeling capabilities and highlight VideoAds as a potentially pivotal benchmark for future research in understanding video that requires high FPS sampling. The dataset and evaluation code will be publicly available at https://videoadsbenchmark.netlify.app.
@@ 1 Introduction
To this end, we introduce VideoAds, the first dedicated dataset for evaluating MLLMs in the context of advertisement video understanding. VideoAds presents unique challenges for video understanding due to its complex temporal structures and high-density semantic content, making it a testbed highly potential for identifying limitations of current MLLMs in temporal reasoning and, accordingly, driving the development of next-generation AI systems with advanced temporal modeling, cross-modal interactions, and multi-stage reasoning capabilities. Beyond academic impact, VideoAds addresses real-world challenges in automated content analysis, brand sentiment prediction, and strategic ad placement, facilitating AI-empowered applications in the advertisement sector.
@@
  • Comprehensive Benchmarking of State-of-the-Art Models: We evaluate leading MLLMs, including GPT-4o [14], Gemini 1.5 Pro [32], and opensource models like LLaVa-Video [48], Qwen2.5-VL [34] over multiple-choice VQA tasks, revealing significant gap from human performance when the models attempt to reason across complex temporal structures. Moreover, MLLMs perform well in finding visual patterns from videos but struggle with video summarization and reasoning, while the opposite is true for humans.
  • Insights about Audio and Chain of Thought in MLLMs: We conduct further analyses about the influence of speech audio transcripts and Chain of Thoughts (CoT). We find that big MLLM models benefit from CoT more than small models, aligning with LMMs’ properties, and the additional information from speech transcripts almost always improves MLLMs’ performance. These results signify the need for future MLLMs to reinforce cross-modal interactions and reasoning.
@@ 2 What makes an advertisement video?
Advertisement videos are a unique form of visual storytelling, meticulously crafted to capture audience attention, convey persuasive messages, and influence consumer behavior within a short time [15, 30]. Unlike traditional video content, advertisements are strategically designed to create brand awareness and drive sales. Their distinctive narrative structure and high-density information make them exceptionally valuable. Consequently, advertisement videos are significantly more expensive than other video types due to high production quality, professional cast, and strategic creative direction.
For example, the second video in Figure 2, “CHANEL N°5 the Film” is a 120-second short advertisement directed by Baz Luhrmann and with a budget of \$33 million  [38]. This advertisement tells a story: “A famous celebrity runs away in a pink dress in the middle of Times Square in New York City, only to get into a cab with a man who does not recognize her. After four days in his Lower East Side apartment, her secretary insists her to return to her life as a celebrity.” This advertisement video shares a plotline similar to Roman Holiday but condenses it into just 120 seconds and provides us with high-density information that has rarely been seen in other video types. The high commercial value translates into complex narrative structures and information-dense sequences, making advertisement videos more challenging to analyze for the current MLLMs which often rely on low FPS sampling.
@@ 3 VideoAds
To build the VideoAds benchmark, we take three steps, video collection followed by manual filtering, curation of VQA tasks, and finally, quality review.
@@ 3.1 Video collection and filtering
We begin by collecting YouTube advertisement videos, using targeted keyword searches to encompass various commercial domains. Specifically, we retrieve videos using commercial search terms about Cars, Sports, Movie Trailers, Food, Tech, Health, Travel, and Financial Services. This broad list ensured our dataset can cover multiple industries, allowing for diverse visual storytelling techniques, product promotions, and narrative-driven ad structures. After this step, we collected 2,700 videos from YouTube.
@@ 3.2 VQA annotations
  • Visual finding: This category assesses the model’s ability to extract concrete visual patterns from video frames. It includes but is not limited to object recognition, attribute identification (e.g., color, shape, size), detecting spatial relationships among objects, and scene recognition. Some exemplar questions are as follows. “What color is the second car in the advertisement?” “Where is the ball on the table?” “Which city does the first scene take place?”
  • Visual summary: This category evaluates the model’s ability to summarize key events and thematic elements in the video. Questions focus on event description, sequential understanding, changes across time, and topic discovery. For example, “How many types of insurance are mentioned in the ads?”, “What is the main product being advertised?”, “What happened before the targted product is introduced?”
@@
While semi-automated VQA annotation methods using powerful MLLMs [42, 20] can easily generate a large number of question-answer pairs using human-provided captions, they are inherently limited by the granularity of the captions and MLLMs’ potential bias.Hence, in this work, the VQA tasks are mainly generated by human annotations. Human annotators are required to create a question, a correct answer, and an incorrect (yet confusing) answer for each VQA task. For every video, an annotator is required to generate 10 VQA pairs, leading to 2,000 VQA tasks in total. We then employ the LLava-Next-72B model [48] to generate four candidate wrong answers. Then, OpenAI GPT4o [13] is used to merge the question, correct answer, and human-provided and model-generated wrong answers into standardized four-option multiple-choice questions. We show examples of all stages of this annotation process in the Appendix.
@@ 3.3 Quality review
To ensure the accuracy and depth of the generated VQA tasks, we apply a quality control step to preserve the most informative and challenging questions, ensuring a balanced distribution across the three primary VQA categories and removing trivial queries. Specifically, we remove VQA pairs that are trivial or overly simplistic, as well as those that contain ambiguities, multiple correct answers, or annotation errors. After expert filtering, we retain 1,100 high-quality VQA tasks from the original 2,000. The numbers of visual finding, visual summary, and visual reasoning tasks are 425, 312, and 363, respectively, and evenly distributed across all question types. Overall, this refinement significantly improved the VQA quality, making it a challenging benchmark for assessing MLLMs.
@@ 3.4.1 Definition of video complexity
Given a specific video $V$, we sample one frame $I_{i}$ per second from it and denote by $n$ the total video frames. Denote by $[-d,d]$ the neighbor of any frame. We define complexity duration $D_{cpx}$ as:
$D_{cpx}=n-\sum_{i=1}^{n}\frac{\sum_{j=\max(i-d,1)}^{\min(i+d,n)}e^{-\frac{|j-i|}{2d}}\cos(f(I_{i}),f(I_{j}))}{\sum_{j=\max(i-d,1)}^{\min(i+d,n)}e^{-\frac{|j-i|}{2d}}},$
@@ 3.4.2 VideoAds analysis and comparison with others
VideoAds distinguishes itself from existing video benchmarks by presenting a unique combination of high video complexity and challenging question-answer (QA) pairs, despite containing relatively short videos: videos in [9] are up to hours. Unlike conventional datasets, such as MSVD-QA[41], Youcook2 [50], and ActivityNet-QA [44], which primarily focus on action recognition or object identification with lower video complexity, VideoAds emphasizes complex temporal reasoning on challenging tasks. This is achieved by curating advertisement videos that compress rich visual narratives and multi-stage events into one to two minutes, demanding long-range contextual tracking and multi-hop reasoning. Compared to Video-MME [8], MMBench-Video [18], and recent Video-MMMU [12], which focus on longer video durations or multi-domain understanding, VideoAds presents a higher density of temporal changes. Table 1 uses the video complexity measure to highlight that VideoAds videos contain significantly more temporal dynamics than other benchmarks. This compact complexity makes it particularly challenging for MLLMs to maintain temporal coherence and causal reasoning, even more so than processing long-form but slowly progressing videos. Furthermore, unlike datasets [18, 9] that use free-form questions, VideoAds employs four-option multiple-choice questions facilitating easy evaluations across different models. Finally, Table 2 presents some statistics for each type of question. In summary, VideoAds raises the bar in video understanding by introducing complex and diverse advertisement videos and VQA tasks.
@@ 4 Experiments and Results
Table 3: Benchmark results for different MLLMs. We can observe that human performance substantially surpasses all SOTA MLLMs while Qwen2.5-VL (72B) leads the best performance of 73.35% under 128 sampling frames. Some prediction cases are shown in the Appendix.
@@ 4.1 Experiment setting
MLLMs: We evaluate our collected dataset via open-source MLMMs including InternVideo2 [35], LongVA [47], InternVL2 [7], LLaVA-Next-Video [48], LLaVA-Video [49], Qwen2.5-VL [2], which covering models from 7B to 72B parameters; and commercial models GPT-4o [14] and Gemini 1.5 Pro [29]. We also compare the influence of the number of uniformly sampled frames on various models for those models that accept different numbers of frames as input.
LLM performance: To ensure VideoAds serves as a reliable benchmark for video understanding, it is crucial to eliminate the possibility of MLLMs exploiting textual patterns rather than genuinely understanding visual content. Therefore, we conduct a baseline experiment using GPT-4 turbo [1] with text-only input, excluding all visual signals. By establishing this text-only baseline, we quantify the extent of textual bias and ensure that VideoAds accurately evaluates temporal reasoning and multi-modal comprehension.
Human performance: To assess the performance of humans in such a challenging task, we recruit Master students as human experts and instruct them to complete the following tests: Each participant is instructed to watch each video once and then answer the corresponding VQA pairs without revisiting the video. This approach mimics the real-world constraints that MLLMs face, ensuring a fair comparison between humans’ and models’ performance.
Evaluation: To ensure a standardized and fair comparison across different MLLMs, we organize the VideoAds dataset into a four-option multiple-choice format. Each question is accompanied by one correct answer and three carefully curated distractors, which are semantically plausible but incorrect. We merge our datasets into the LMM-Eval package [46], which has been widely used in MLLMs evaluation such that other researchers can use this dataset easily.
@@ 4.2 Benchmarking results
Current MLLMs excel in static visual recognition but struggle with complex reasoning, while humans give contradictory results: From Table 3, a clear performance disparity emerges across the three task types: visual finding, visual summary, and visual reasoning. Specifically, most MLLMs perform the best on the visual finding tasks, followed by visual summary, with the worst performance observed on Visual Reasoning tasks. For example, GPT-4o achieves an accuracy of 75.06% on the visual finding tasks but drops significantly to 62.26% on the visual reasoning tasks with 128 frames. This pattern is consistent across other models, such as LLaVA-Video and Qwen2.5-VL under different sampling frames. This performance gradient suggests that current MLLMs are more adept at static visual recognition and shallow temporal tracking but struggle with tasks requiring complex narrative reasoning. We also can observe an interesting result for human performance in Table 3 that the performance on the most challenging reasoning tasks (accuracy 95.04%) is even higher than the visual finding (accuracy 93.41%). This is reasonable given that human experts find it easier to find the visual reasoning structures in advertisement videos at a high level but find it hard to remember every detail.
State-of-the-art opensource MLLMs can beat GPT-4o and Gemini 1.5 Pro: For a long time commercial LLMs [23, 14, 32] can easily beat all opensource models in video understanding while VideoAds lead to the opposite result with Qwen2.5-VL outperforming GPT-4o and Gemini 1.5 Pro in Table 3. GPT-4o and Gemini 1.5 Pro can still achieve SOTA results (above 75%) on the visual finding task but struggle with the visual summary and reasoning (around 66%-67%). This further emphasizes that rather than taking video as a grid of images, it is necessary to train MLLMs on video data specifically to extract the temporal relationship within the video.
Influence of increasing sampled frames varies across the models: When a video is simple, a limited set of sampled frames can include all necessary information. However, in complex videos like those in VideoAds, increasing the number of frames can significantly impact performance. We observe that Qwen2.5-VL benefits more from additional frames than GPT-4o and LLaVA-Video. The latter two models show notable improvements in visual finding tasks but only marginal gains in reasoning tasks when we fed them frames per video from 32 to 128. We hypothesize due to high video complexity in VideoAds, it actually requires the MLLMs to process high fps and long context tokens, which are underexplored in the previous work [34, 29, 14, 48].

### v2

@@ Abstract
Advertisement videos serve as a rich and valuable source of purpose-driven information, encompassing high-quality visual, textual, and contextual cues designed to engage viewers. They are often more complex than general videos of similar duration due to their structured narratives and rapid scene transitions, posing significant challenges to multi-modal large language models (MLLMs). In this work, we introduce VideoAds, the first dataset tailored for benchmarking the performance of MLLMs on advertisement videos. VideoAds comprises well-curated advertisement videos with complex temporal structures, accompanied by manually annotated diverse questions across three core tasks: visual finding, video summary, and visual reasoning. We propose a quantitative measure to compare VideoAds against existing benchmarks in terms of video complexity. Through extensive experiments, we find that Qwen2.5-VL-72B, an open‑source MLLM, achieves 73.35% accuracy on VideoAds, outperforms GPT-4o (66.82%) and Gemini-1.5 Pro (69.66%); the two proprietary models especially fall behind the open‑source model in video summarization and reasoning, but perform the best in visual finding. Gemini-2.5 Pro leads with an accuracy of 80.04%. Notably, human experts easily achieve a remarkable accuracy of 94.27%. These results underscore the necessity of advancing MLLMs’ temporal modeling capabilities and highlight VideoAds as a potentially pivotal benchmark for future research in understanding video that requires high FPS sampling. The dataset and evaluation code will be publicly available at https://videoadsbenchmark.netlify.app.
@@ 1 Introduction
To this end, we introduce VideoAds, the first dedicated dataset for evaluating MLLMs in the context of advertisement video understanding. It is important to position VideoAds in the landscape of existing benchmarks. VideoAds is not intended to replace larger, more comprehensive benchmarks like Video-MME [7], but rather to complement them. VideoAds presents unique challenges for video understanding due to its complex temporal structures and high-density semantic content, making it a testbed with high potential for identifying limitations of current MLLMs in temporal reasoning and, accordingly, driving the development of next-generation AI systems with advanced temporal modeling, cross-modal interactions, and multi-stage reasoning capabilities. Beyond academic impact, VideoAds addresses real-world challenges in automated content analysis, brand sentiment prediction, and strategic ad placement, facilitating AI-empowered applications in the advertisement sector.
@@
  • Comprehensive Benchmarking of State-of-the-Art Models: We evaluate leading MLLMs, including GPT-4o [12], Gemini 1.5 Pro [31], and open source models like LLaVa-Video [48], Qwen2.5-VL [33] over multiple-choice VQA tasks, revealing significant gap from human performance when the models attempt to reason across complex temporal structures. Moreover, MLLMs perform well in finding visual patterns from videos but struggle with video summarization and reasoning.
  • Insights about Audio and Chain of Thought in MLLMs: We conduct further analyses about the influence of speech audio transcripts and Chain of Thoughts (CoT). We find that large MLLMs benefit from CoT more than small models, aligning with LMMs’ properties, and the additional information from speech transcripts almost always improves MLLMs’ performance. These results signify the need for future MLLMs to reinforce cross-modal interactions and reasoning.
@@ 2 What makes an advertisement video?
Advertisement videos are a unique form of visual storytelling, meticulously crafted to capture audience attention, convey persuasive messages, and influence consumer behavior within a short time [13, 29]. For example, the second video in Figure 2, “CHANEL N°5 the Film” is a 120-second short advertisement directed by Baz Luhrmann and with a budget of \$33 million [37]. This advertisement tells a story: “A famous celebrity runs away in a pink dress in the middle of Times Square in New York City, only to get into a cab with a man who does not recognize her. After four days in his Lower East Side apartment, her secretary insists that she return as a celebrity.” This advertisement video shares a plotline similar to Roman Holiday but condenses it into just 120 seconds and provides us with high-density information that has rarely been seen in other video types. The high commercial value translates into complex narrative structures and information-dense sequences, making advertisement videos more challenging to analyze for the current MLLMs which often rely on low FPS sampling.
@@ 3 VideoAds
To build the VideoAds benchmark, we take three steps: video collection followed by manual filtering; curation of VQA tasks; and quality review.
@@ 3.1 Video collection and filtering
We begin by collecting YouTube advertisement videos, using targeted keyword searches to encompass various commercial domains. Specifically, we retrieve videos using commercial search terms about Cars, Sports, Movie Trailers, Food, Tech, Health, Travel, and Financial Services. This broad list ensures our dataset covers multiple industries, allowing for diverse visual storytelling techniques, product promotions, and narrative-driven ad structures. After this step, we collected 2,700 videos from YouTube.
@@ 3.2 VQA annotations
  • Visual finding: This category assesses the model’s ability to extract concrete visual patterns from video frames. It includes but is not limited to object recognition, attribute identification (e.g., color, shape, size), detecting spatial relationships among objects, and scene recognition. Some exemplar questions are as follows. “What color is the second car in the advertisement?” “Where is the ball on the table?” “In which city does the first scene take place?”
  • Visual summary: This category evaluates the model’s ability to summarize key events and thematic elements in the video. Questions focus on event description, sequential understanding, changes across time, and topic discovery. For example, “How many types of insurance are mentioned in the ads?”, “What is the main product being advertised?”, “What happened before the targeted product is introduced?”
@@
While semi-automated VQA annotation methods using powerful MLLMs [41, 19] can easily generate a large number of question-answer pairs using human-provided captions, they are inherently limited by the granularity of the captions and MLLMs’ potential bias. Hence, in this work, the VQA tasks are mainly generated by human annotations. Human annotators are required to create a question, a correct answer, and an incorrect (yet confusing) answer for each VQA task. For every video, an annotator is required to generate 10 VQA pairs, leading to 2,000 VQA tasks in total. We then employ the LLava-Next-72B model [48] to generate four candidate incorrect answers. Then, OpenAI GPT4o [11] is used to merge the question, correct answer, and human-provided and model-generated incorrect answers into standardized four-option multiple-choice questions. We show examples of all stages of this annotation process in the Appendix.
@@ 3.3 Quality review
To ensure the accuracy and depth of the generated VQA tasks, we apply a quality control step to preserve the most informative and challenging questions, ensuring a balanced distribution across the three primary VQA categories and removing trivial queries. Specifically, we remove VQA pairs that are trivial or overly simplistic, as well as those that contain ambiguities, multiple correct answers, or annotation errors. After expert filtering, we retain 1,100 high-quality VQA tasks from the original 2,000. The numbers of visual finding, visual summary, and visual reasoning tasks are 425, 312 and 363, respectively, and evenly distributed across all question types. Overall, this refinement significantly improved the VQA quality, making it a challenging benchmark for assessing MLLMs.
@@ 3.4.1 Definition of video complexity
Given a specific video $V$, we sample one frame $I_{i}$ per second from it and denote by $n$ the total number of video frames. Denote by [d,d][-d,d] the neighbor of any frame. We define complexity duration $D_{cpx}$ as:
$D_{cpx}=n-\sum_{i=1}^{n}\frac{\sum_{j=\max(i-d,1)}^{\min(i+d,n)}e^{-\frac{|j-i|}{2d}}\cos(f(I_{i}),f(I_{j}))}{\sum_{j=\max(i-d,1)}^{\min(i+d,n)}e^{-\frac{|j-i|}{2d}}},$ (1)
@@ 3.4.2 VideoAds analysis and comparison with others
VideoAds distinguishes itself from existing video benchmarks by presenting a unique combination of high video complexity and challenging question-answer (QA) pairs, despite containing relatively short videos: videos in [8] are up to hours long. Unlike conventional datasets, such as MSVD-QA [40], Youcook2 [50], and ActivityNet-QA [44], which primarily focus on action recognition or object identification with lower video complexity, VideoAds emphasizes complex temporal reasoning on challenging tasks. This is achieved by curating advertisement videos that compress rich visual narratives and multi-stage events into one to two minutes, demanding long-range contextual tracking and multi-hop reasoning. Compared to Video-MME [7], MMBench-Video [17], and recent Video-MMMU [10], which focus on longer video durations or multi-domain understanding, VideoAds presents a higher density of temporal changes. Table 1 uses the video complexity measure to highlight that VideoAds videos contain significantly more temporal dynamics than other benchmarks. This compact complexity makes it particularly challenging for MLLMs to maintain temporal coherence and causal reasoning, even more so than processing long-form but slowly progressing videos. Furthermore, unlike datasets [17, 8] that use free-form questions, VideoAds employs four-option multiple-choice questions facilitating easy evaluations across different models. Finally, Table 2 presents some statistics for each type of question. In summary, VideoAds raises the bar in video understanding by introducing complex and diverse advertisement videos and VQA tasks.
It is crucial to understand why benchmarks like Temp-Compass and MVBench score lower on our complexity metric despite being used for fine-grained temporal understanding. Our metric, based on DINO-v2 feature variance, excels at capturing coarse visual shifts and significant content changes typical of advertisements with rapid cuts between distinct scenes. In contrast, videos in TempCompass or MVBench often feature more subtle temporal changes (e.g., an object moving slightly), which results in simpler overall video content and thus a lower complexity score. VideoAds therefore, serves as a complementary benchmark, focused on challenging models with fast-paced, semantics-rich visual transitions rather than fine-grained temporal shifts.
@@ 4 Experiments and Results
Table 3: Benchmark results for different MLLMs. We can observe that human performance substantially surpasses all SOTA MLLMs, while Gemini 2.5 Pro leads the best performance of 80.04%. Some prediction cases are shown in the Appendix.
@@ 4.1 Experiment setting
MLLMs: We evaluate our collected dataset using open-source MLMMs, including InternVideo2 [34], LongVA [47], InternVL2 [5], LLaVA-Next-Video [48], LLaVA-Video [49], Qwen2.5-VL [1], which covers models from 7B to 72B parameters; and commercial models GPT-4o [12] , Gemini 1.5 Pro and 2.5 Pro [28, 6]. We also compare the influence of the number of sampled frames on various models for those models that accept different numbers of frames as input.
LLM performance: To ensure VideoAds serves as a reliable benchmark for video understanding, it is crucial to eliminate the possibility of MLLMs exploiting textual patterns rather than genuinely understanding visual content. Therefore, we conduct a baseline experiment using GPT-4o with text-only input, excluding all visual signals. Specifically, the model was prompted with instructions such as, ’Based only on the provided question, answer the following multiple-choice question,’ without any indication that visual input was required. This process helps quantify the model’s reliance on visual information versus textual bias.
Human performance: To assess the performance of humans in such a challenging task, we recruit two master’s students as human experts and instruct them to complete the following tests: Given that most advertisement videos are designed to contain specific information for human, it is relatively easy to answer accurately by the evaluators. Each participant is instructed to watch each video once and then answer the corresponding VQA pairs without revisiting the video. This approach mimics the real-world constraints that MLLMs face, ensuring a fair comparison between human and model performance.
Evaluation: To ensure a standardized and fair comparison across different MLLMs, we organize the VideoAds dataset into a four-option multiple-choice format. Each question is accompanied by one correct answer and three carefully curated distractors, which are semantically plausible but incorrect. We merge our dataset into the LMM-Eval package [46], which has been widely used in MLLMs evaluation such that other researchers can use this dataset easily.
@@ 4.2 Benchmarking results
Current MLLMs excel in static visual recognition but struggle with complex reasoning, while humans give contradictory results: From Table 3, a clear performance disparity emerges across the three task types: visual finding, visual summary, and visual reasoning. Specifically, most MLLMs perform the best on the visual finding tasks, followed by visual summary, with the worst performance observed on Visual Reasoning tasks. For example, GPT-4o achieves an accuracy of 75.06% on the visual finding tasks but drops significantly to 62.26% on the visual reasoning tasks with 128 frames. This pattern is consistent across other models, such as LLaVA-Video and Qwen2.5-VL under different sampling frames. This performance gradient suggests that current MLLMs are more adept at static visual recognition and shallow temporal tracking but struggle with tasks requiring complex narrative reasoning. We also observe an interesting result for human performance in Table 3 that the performance on the most challenging reasoning tasks (accuracy 95.04%) is even higher than the visual finding (accuracy 93.41%). This is reasonable given that human experts find it easier to find the visual reasoning structures in advertisement videos at a high level but find it hard to remember every detail.
State-of-the-art open source MLLMs can beat GPT-4o and Gemini 1.5 Pro, while Gemini 2.5 Pro continues to lead performance: For a long time commercial LLMs [22, 12, 31] can easily beat all opensource models in video understanding. VideoAds, however, produces the opposite result, with Qwen2.5-VL outperforming GPT-4o and Gemini 1.5 Pro in Table 3. GPT-4o and Gemini 1.5 Pro can still achieve SOTA results (above 75%) on the visual finding task but struggle with the visual summary and reasoning (around 66%-67%). However, Gemini 2.5 Pro, the most recent model, keep leading the performance with an accuracy of 80.04%. This further emphasizes that rather than taking video as a grid of images, it is necessary to train MLLMs on video data specifically to extract the temporal relationship within the video like used in Gemini 2.5 Pro [6].
Influence of increasing sampled frames varies across the models: When a video is simple, a limited set of sampled frames can include all necessary information. However, in complex videos like those in VideoAds, increasing the number of frames can significantly impact performance. We observe that Qwen2.5-VL benefits more from additional frames than GPT-4o and LLaVA-Video. The latter two models show notable improvements in visual finding tasks but only marginal gains in reasoning tasks when we increased the number of frames per video from 32 to 128. We hypothesize that, due to high video complexity in VideoAds, it actually requires the MLLMs to process high fps and long context tokens, which are underexplored in the previous work [33, 28, 12, 48].